6. If the lateral area of a right circular cylinder is 90 cm2 and its volume is 225 cm3, find its radius.
1
Expert's answer
2020-12-10T20:23:23-0500

Consider a right circular cylinder with radius of base "r" and height "h"



The volume of the cylinder is "V=\pi r^2h."

The lateral area of the cylinder is "F=2\pi rh"

Then


"\dfrac{V}{F}=\dfrac{\pi r^2 h}{2\pi rh}=\dfrac{r}{2}"

"r=\dfrac{2V}{F}"

Given "F=90\ cm^2, V=225\ cm^3." Then


"r=\dfrac{2(225\ cm^3)}{90\ cm^2}=5 \ cm"

The radius of the cylindeer is 5 cm.